A plastic rod has been charged to -30 nC by rubbing.
a. Have electrons been added to the rod or electrons removed? Explain.b. How many electrons or protons have been removed?
Answer #1

(a) As the rod attains a net negative charge, this means excess of electrons. Hence electrons are added to the rod.

(b) Now charge on an electron = 1.6*10^-19 C

Charge on the rod = -30 nC = -30*10^-9 C

Let x number of electrons has been added, then,

x*(-1.6*10^-19) = - 30*10^-9

=> x = -30*10^-9/-1.6*10^-19 = 1875*10^8

So, number of electrons added = 1875*10^8